A particle moving along the x axis . It’s position varies with time according to x(t)=-4t+2t^2

a) Find the average velocity between t=0s and t=2s

b) Find the average speed of the particle between t=0s and t=2s


1
Expert's answer
2021-11-07T19:27:44-0500

a)

average velocity:

"v=\frac{x(2)-x(0)}{2-0}=\frac{0}{2}=0"


b)

distance travelled = length of curve x(t):


"L=\int^2_0\sqrt{1+(x'(t))^2}dt"


"x'(t)=4t-4"


"L=\int^2_0\sqrt{1+(4t-4)^2}dt=(\frac{arcsinh(4t-4)+(4t-4)\sqrt{16t^2-32t+17}}{8})|^2_0="


"=2.32+2.32=4.64"


average speed:

"v=\frac{L}{\Delta t}=\frac{4.64}{2}=2.32" m/s